OpenStudy (anonymous):

-4(2n-5)=-8n-20 step by step

OpenStudy (anonymous):

need help please

OpenStudy (katrinakaif):

Hmm n turns out to be 0..

OpenStudy (anonymous):

can you show me the steps to figure out the equation please

OpenStudy (katrinakaif):

Yeah sure: -4(2n - 5) = -8n - 20 Distribute -4 -8n + 20 = -8n -20 -8n + 8n = -20 - 20 0n = -40 n = 0